FIND THE GENERAL SOLUTION

Need ALL 4. Differential Equations class

13. x(t) E) (1) 2) st) 14. 7(t) = ( 1 5 0 15. 7(t) = ( 2 1 32 0 -20 x(t) - 1 16. 7(t) = 3 0 2 1 -10 | z(t) -2 -1
Answer #1

332: Pintos font tirint) eth (2 cost-sort to Jeneral solution. 9/4= Get truse- 기 it) til nt Ducato d=288, A153) 3( mt JU -5-9

geneuassol! alt) = ce 2 331 (1-a renn [sinst the corre + Che cost sunt 2005/3t et a 251047- 16 A o 2 O = -2, - 12/21 3-| A -2